A consumer spend income 5000 on food and cloth draw the budget line and identify intercept food price is100 and cloth price is 250

1
Expert's answer
2022-01-23T15:37:45-0500

Question #289904

Given:


"P_x = 100"


"P_y = 250"


"Budget=5,000"

Solution:

Let "x" be the quantity of food consumed by the customer, "y" be the quantity of cloth consumed by the customer. In this case, the budget line has the equation:

"5,000 = P_x \\cdot x + P_y\\cdot y \t\\Rightarrow"

"\Rightarrow 5,000=100\cdot x+250\cdot y"

The intercepts are:

  • For the x-intercept, we have:
"y=0\t\\Rightarrow5,000=100\\cdot x+250\\cdot0\t\\Rightarrow"


"\Rightarrow 5,000=100\cdot x\Rightarrow x=5,000\div 100=50"


Thus, x-intercept is (50, 0).

  • For the y-intercept, we have:
"x=0\Rightarrow 5,000=100\cdot 0+250 \cdot y\Rightarrow"

"\Rightarrow 5,000 = 250\cdot y\Rightarrow y=5,000\div 250=20"

Thus, the y-intercept is (0,20).

Based on the equation of the budget line, the graph is:


Answer: the equation of the budget line is "5,000=100\cdot x+250\cdot y" . The intercepts are: (0, 20) and (50,0).